# Evaluating Named Entity Recognition: Comparative Analysis of Mono- and Multilingual Transformer Models on Brazilian Corporate Earnings Call Transcriptions

### v1

@@ 9 Final Considerations
Upon conducting manual error analysis, we identified issues stemming from the annotation process using weak supervision techniques. Although regular expressions and heuristics are useful the annotation, they may not cover all possible cases, particularly in instances involving money units such as millions and billions, where annotations were, in some cases, limited to currency symbols and numerical values only. Additionally, we observed challenges in annotating nested entities or multi-word entities, as well as variations in entity spellings or formats, such as differences between singular and plural forms. These limitations are inherent to weak supervision approaches. Nonetheless, despite these challenges, the dataset remains a valuable resource for further studies in the field.
@@
Our approach to framing the token classification problem as a text generation task yielded promising results, notably with T5 models generating outputs closely resembling the original sentences. However, we encountered critical issues where the model altered monetary and percentage values, which are crucial in the financial domain. Despite these challenges, our strategy proved effective, as we fine-tuned models with fewer parameters and achieved higher F1-scores compared to related work Wang2022instructionner .
As part of future work, we plan to expand the dataset by incorporating additional transcriptions, given that companies typically release them quarterly. Addressing weaknesses in weak supervision techniques will be a priority to enhance dataset quality, alongside a revision of the labels list to ensure its comprehensiveness and accuracy. Moreover, we aim to conduct experiments aimed at determining the optimal number of entities required to achieve higher performance and error metrics in NER models. This approach will facilitate human annotation efforts and enable more efficient utilization of computational resources. Furthermore, we aim to explore the capabilities of generative models in addressing critical issues such as generating accurate monetary and percentage values and preventing the repetition of sentence parts.
@@ Appendix A Results of inference using Test dataset per class and model
Precision Recall F1-score Support
BALANCO_PATRIMONIAL 0.9915 0.9832 0.9873 119
CARTEIRA 0.9996 0.9968 0.9982 2836
CLIENTE 1.0000 0.9993 0.9996 1387
COMPANY 0.9927 0.9976 0.9952 1647
CONDICOES_MACROECONOMICAS 0.9980 0.9941 0.9961 511
DESPESA 1.0000 1.0000 1.0000 703
INDICADOR_EFICIENCIA 1.0000 1.0000 1.0000 184
INDICADOR_LIQUIDEZ 0.9978 1.0000 0.9989 458
INDICADOR_RENTABILIDADE 0.9799 0.9966 0.9882 294
INDICADOR_VALUATION 1.0000 0.9452 0.9718 73
LUCRO 1.0000 1.0000 1.0000 135
MONEY 0.9975 0.9896 0.9935 1629
ORG 1.0000 1.0000 1.0000 91
PERCENTUAL 0.9978 0.9943 0.9961 5092
PRODUTO 1.0000 0.9902 0.9951 511
PROVENTO 1.0000 1.0000 1.0000 110
PROVISAO 0.9980 1.0000 0.9990 1479
QUARTER 1.0000 1.0000 1.0000 424
RECEITA 0.9883 0.9883 0.9883 171
RESULTADO 0.9286 0.8125 0.8667 16
RISCO 0.9991 0.9974 0.9983 2341
SEMESTER 1.0000 0.9941 0.9971 170
YEAR 0.9949 1.0000 0.9975 392
micro avg 0.9977 0.9958 0.9968 20773
macro avg 0.9941 0.9861 0.9899 20773
weighted avg 0.9977 0.9958 0.9968 20773
@@
Precision Recall F1-score Support
BALANCO_PATRIMONIAL 1.0000 0.9788 0.9893 283
CARTEIRA 0.9996 0.9976 0.9986 2863
CLIENTE 0.9986 0.9986 0.9986 1405
COMPANY 0.9918 0.9975 0.9946 1573
CONDICOES_MACROECONOMICAS 0.9987 0.9962 0.9975 797
DESPESA 0.9994 1.0000 0.9997 1580
INDICADOR_EFICIENCIA 1.0000 0.9915 0.9957 353
INDICADOR_LIQUIDEZ 0.9902 0.9951 0.9926 406
INDICADOR_RENTABILIDADE 0.9933 0.9900 0.9917 300
INDICADOR_VALUATION 1.0000 0.9600 0.9796 50
LUCRO 1.0000 0.9921 0.9960 252
MONEY 0.9963 0.9902 0.9932 1626
ORG 0.9890 0.9890 0.9890 91
PERCENTUAL 0.9986 0.9948 0.9967 5046
PRODUTO 1.0000 0.9850 0.9924 933
PROVENTO 1.0000 1.0000 1.0000 121
PROVISAO 1.0000 0.9986 0.9993 1465
QUARTER 0.9930 0.9977 0.9953 427
RECEITA 1.0000 1.0000 1.0000 497
RESULTADO 0.7857 0.9167 0.8462 12
RISCO 0.9975 0.9975 0.9975 1985
SEMESTER 1.0000 1.0000 1.0000 169
YEAR 0.9969 0.9938 0.9954 325
micro avg 0.9978 0.9955 0.9966 22559
macro avg 0.9882 0.9896 0.9886 22559
weighted avg 0.9978 0.9955 0.9966 22559
@@
Precision Recall F1-score Support
BALANCO_PATRIMONIAL 1.0000 0.9833 0.9916 120
CARTEIRA 0.9930 0.9916 0.9923 1431
CLIENTE 0.9985 0.9977 0.9981 1292
COMPANY 0.9851 0.9960 0.9905 1262
CONDICOES_MACROECONOMICAS 0.9903 0.9903 0.9903 308
DESPESA 0.9981 0.9981 0.9981 527
INDICADOR_EFICIENCIA 1.0000 1.0000 1.0000 173
INDICADOR_LIQUIDEZ 0.9749 0.9831 0.9790 237
INDICADOR_RENTABILIDADE 0.9688 0.9894 0.9789 94
INDICADOR_VALUATION 1.0000 0.9231 0.9600 26
LUCRO 1.0000 0.9921 0.9960 126
MONEY 0.9889 0.9828 0.9858 814
ORG 0.9780 1.0000 0.9889 89
PERCENTUAL 0.9943 0.9943 0.9943 2629
PRODUTO 0.9881 0.9811 0.9846 423
PROVENTO 1.0000 0.9836 0.9917 61
PROVISAO 0.9985 0.9970 0.9977 656
QUARTER 0.9911 0.9911 0.9911 336
RECEITA 0.9641 0.9641 0.9641 167
RESULTADO 0.9167 1.0000 0.9565 11
RISCO 0.9986 0.9946 0.9966 735
SEMESTER 0.9910 0.9735 0.9821 113
YEAR 0.9175 0.9858 0.9504 282
micro avg 0.9905 0.9921 0.9913 11912
macro avg 0.9841 0.9866 0.9852 11912
weighted avg 0.9907 0.9921 0.9914 11912
@@
Precision Recall F1-score Support
BALANCO_PATRIMONIAL 1.0000 0.9916 0.9958 119
CARTEIRA 0.9979 0.9972 0.9976 1430
CLIENTE 0.9985 0.9977 0.9981 1292
COMPANY 0.9890 0.9992 0.9941 1263
CONDICOES_MACROECONOMICAS 0.9968 0.9968 0.9968 308
DESPESA 0.9943 0.9943 0.9943 527
INDICADOR_EFICIENCIA 1.0000 1.0000 1.0000 173
INDICADOR_LIQUIDEZ 0.9958 0.9958 0.9958 239
INDICADOR_RENTABILIDADE 0.9792 0.9895 0.9843 95
INDICADOR_VALUATION 1.0000 1.0000 1.0000 24
LUCRO 1.0000 0.9921 0.9960 126
MONEY 0.9827 0.9719 0.9773 818
ORG 0.9890 1.0000 0.9945 90
PERCENTUAL 0.9951 0.9883 0.9917 2647
PRODUTO 1.0000 0.9859 0.9929 426
PROVENTO 0.9833 0.9833 0.9833 60
PROVISAO 0.9939 0.9939 0.9939 655
QUARTER 0.9940 0.9911 0.9926 337
RECEITA 1.0000 1.0000 1.0000 167
RESULTADO 0.9167 1.0000 0.9565 11
RISCO 0.9959 0.9959 0.9959 732
SEMESTER 1.0000 0.9823 0.9911 113
YEAR 0.9868 0.8519 0.9144 351
micro avg 0.9942 0.9883 0.9912 12003
macro avg 0.9908 0.9869 0.9885 12003
weighted avg 0.9942 0.9883 0.9911 12003

### v2

@@ 7 Conclusions
Through manual error analysis, we identified issues from the weakly supervised annotation process. While regular expressions and heuristics are helpful, they don’t cover all cases, particularly with money units like millions and billions, where annotations were sometimes limited to currency symbols and numbers. We also observed difficulties with nested or multi-word entities and variations in entity spellings or formats, such as singular and plural forms. These limitations are inherent to weak supervision. Despite these challenges, the dataset remains valuable for further research.
@@
Our approach to framing the token classification problem as a text generation task yielded promising results, notably with T5 models generating outputs closely resembling the original sentences. However, we encountered critical issues where the model altered monetary and percentage values, which are crucial in the financial domain. Despite these challenges, our strategy proved effective, as we fine-tuned models with fewer parameters and achieved higher F1-scores compared to related work [20].
For future work, we plan to expand the dataset with additional transcripts, as companies release them quarterly. Additionally, we aim to improve the dataset quality and revise the labels list for comprehensiveness and accuracy. Furthermore, we intend to explore prompt engineering techniques in generative models.
@@ Appendix A Results of inference using Test dataset per class and model
Precision Recall F1-score
BALANCO_PATRIMONIAL 0.9938 0.9802 0.9867
CARTEIRA 0.9996 0.9968 0.9982
CLIENTE 1.0000 0.9993 0.9996
COMPANY 0.9929 0.9976 0.9952
CONDICOES_MACROECONOMICAS 0.9978 0.9946 0.9962
DESPESA 1.0000 1.0000 1.0000
INDICADOR_EFICIENCIA 1.0000 1.0000 1.0000
INDICADOR_LIQUIDEZ 0.9978 1.0000 0.9989
INDICADOR_RENTABILIDADE 0.9727 0.9943 0.9832
INDICADOR_VALUATION 1.0000 0.9385 0.9636
LUCRO 1.0000 1.0000 1.0000
MONEY 0.9974 0.9892 0.9933
ORG 1.0000 1.0000 1.0000
PERCENTUAL 0.9979 0.9943 0.9961
PRODUTO 1.0000 0.9913 0.9956
PROVENTO 1.0000 1.0000 1.0000
PROVISAO 0.9982 1.0000 0.9991
QUARTER 1.0000 1.0000 1.0000
RECEITA 0.9818 0.9818 0.9818
RESULTADO 0.9500 0.8433 0.8878
RISCO 0.9992 0.9973 0.9982
SEMESTER 1.0000 0.9950 0.9975
YEAR 0.9943 1.0000 0.9971
macro avg 0.9945 0.9867 0.9899
micro avg 0.9977 0.9958 0.9968
weighted avg 0.9978 0.9958 0.9967
@@
Precision Recall F1-score
BALANCO_PATRIMONIAL 1.0000 0.9771 0.9882
CARTEIRA 0.9996 0.9976 0.9986
CLIENTE 0.9985 0.9986 0.9986
COMPANY 0.9920 0.9976 0.9948
CONDICOES_MACROECONOMICAS 0.9987 0.9968 0.9978
DESPESA 0.9994 1.0000 0.9997
INDICADOR_EFICIENCIA 1.0000 0.9930 0.9964
INDICADOR_LIQUIDEZ 0.9901 0.9952 0.9926
INDICADOR_RENTABILIDADE 0.9910 0.9873 0.9891
INDICADOR_VALUATION 1.0000 0.9500 0.9714
LUCRO 1.0000 0.9913 0.9956
MONEY 0.9961 0.9897 0.9929
ORG 0.9923 0.9923 0.9923
PERCENTUAL 0.9986 0.9949 0.9967
PRODUTO 1.0000 0.9861 0.9929
PROVENTO 1.0000 1.0000 1.0000
PROVISAO 1.0000 0.9984 0.9992
QUARTER 0.9931 0.9979 0.9955
RECEITA 1.0000 1.0000 1.0000
RESULTADO 0.8500 0.9000 0.8667
RISCO 0.9976 0.9974 0.9975
SEMESTER 1.0000 1.0000 1.0000
YEAR 0.9968 0.9925 0.9946
macro avg 0.9910 0.9884 0.9892
micro avg 0.9978 0.9955 0.9966
weighted avg 0.9979 0.9955 0.9966
@@
Precision Recall F1-score
BALANCO_PATRIMONIAL 1.0000 0.9802 0.9898
CARTEIRA 0.9929 0.9916 0.9922
CLIENTE 0.9985 0.9977 0.9981
COMPANY 0.9855 0.9962 0.9908
CONDICOES_MACROECONOMICAS 0.9903 0.9910 0.9905
DESPESA 0.9981 0.9981 0.9981
INDICADOR_EFICIENCIA 1.0000 1.0000 1.0000
INDICADOR_LIQUIDEZ 0.9749 0.9831 0.9790
INDICADOR_RENTABILIDADE 0.9549 0.9933 0.9724
INDICADOR_VALUATION 1.0000 0.9278 0.9597
LUCRO 1.0000 0.9913 0.9956
MONEY 0.9887 0.9823 0.9855
ORG 0.9746 1.0000 0.9869
PERCENTUAL 0.9944 0.9943 0.9944
PRODUTO 0.9884 0.9816 0.9850
PROVENTO 1.0000 0.9846 0.9920
PROVISAO 0.9982 0.9968 0.9975
QUARTER 0.9910 0.9910 0.9910
RECEITA 0.9548 0.9548 0.9548
RESULTADO 0.9000 1.0000 0.9333
RISCO 0.9987 0.9948 0.9967
SEMESTER 0.9913 0.9778 0.9838
YEAR 0.9185 0.9853 0.9505
macro avg 0.9823 0.9867 0.9834
micro avg 0.9905 0.9921 0.9913
weighted avg 0.9908 0.9921 0.9914
@@
Precision Recall F1-score
BALANCO_PATRIMONIAL 1.0000 0.9889 0.9943
CARTEIRA 0.9979 0.9972 0.9975
CLIENTE 0.9985 0.9977 0.9981
COMPANY 0.9892 0.9992 0.9942
CONDICOES_MACROECONOMICAS 0.9967 0.9971 0.9969
DESPESA 0.9945 0.9947 0.9946
INDICADOR_EFICIENCIA 1.0000 1.0000 1.0000
INDICADOR_LIQUIDEZ 0.9963 0.9963 0.9963
INDICADOR_RENTABILIDADE 0.9716 0.9867 0.9783
INDICADOR_VALUATION 1.0000 1.0000 1.0000
LUCRO 1.0000 0.9913 0.9956
MONEY 0.9820 0.9706 0.9762
ORG 0.9889 1.0000 0.9943
PERCENTUAL 0.9951 0.9883 0.9917
PRODUTO 1.0000 0.9867 0.9933
PROVENTO 0.9882 0.9882 0.9882
PROVISAO 0.9946 0.9946 0.9946
QUARTER 0.9934 0.9905 0.9919
RECEITA 1.0000 1.0000 1.0000
RESULTADO 0.9000 1.0000 0.9333
RISCO 0.9961 0.9962 0.9961
SEMESTER 1.0000 0.9846 0.9920
YEAR 0.9851 0.9111 0.9353
macro avg 0.9899 0.9896 0.9884
micro avg 0.9942 0.9885 0.9913
weighted avg 0.9943 0.9885 0.9907